OpenStudy (anonymous):

i need help because i keep getting the answer wrong and i dont get it A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Enter your answer in the box. inches Part B How many miles are represented by 1 4 inch on the map? Enter your answer in the box. miles

OpenStudy (igreen):

We can set up a proportion. \(\dfrac{2}{25} = \dfrac{x}{60}\) Multiply 60 to both sides: \(x = \dfrac{2 \times 60}{25}\) SImplify

OpenStudy (igreen):

*Simplify

OpenStudy (igreen):

Multiply 2 * 60, then divide by 25.

OpenStudy (igreen):

Can you do that? @happyb26

OpenStudy (anonymous):

4.8

OpenStudy (igreen):

Yes, that's our answer. 4.8 inches will represent 60 miles.

OpenStudy (igreen):

For Part B, we can set up another proportion. \(\dfrac{2}{25} = \dfrac{0.25}{x}\) Cross multiply: 2x = 25 * 0.25 Multiply 25 * 0.25

OpenStudy (anonymous):

6.25

OpenStudy (anonymous):

can you also help me with another question?

OpenStudy (igreen):

Yes, so we have: 2x = 6.25 Now divide 2 to both sides, what's 6.25 / 2?

OpenStudy (anonymous):

3.125

OpenStudy (igreen):

Yes, that's our answer for Part B.

OpenStudy (igreen):

I can help with another question, but you should close this and open a new one.

OpenStudy (anonymous):

okay

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.